A 20 meter ladder and a 15 meter ladder were leaned against a building. The bottom of the longer ladder was 7 meters farther from the building than the bottom of the shorter ladder, but both ladders reached the same distance up the building. Find the distance the ladders reached up the building.
Answer by richwmiller(17219)

20^2=(x+7)^2+y^2
15^2=x^2+y^2